### v4

# Fast Preemption: Forward-Backward Cascade Learning for Efficient and Transferable Proactive Adversarial Defense

Abstract
Deep learning technology has brought convenience and advanced developments but has become untrustworthy due to its sensitivity to adversarial attacks. Attackers may utilize this sensitivity to manipulate predictions. To defend against such attacks, existing anti-adversarial methods typically counteract adversarial perturbations post-attack, while we have devised a proactive strategy that preempts by safeguarding media upfront, effectively neutralizing potential adversarial effects before the third-party attacks occur. This strategy, dubbed Fast Preemption, provides an efficient transferable preemptive defense by using different models for labeling inputs and learning crucial features. A forward-backward cascade learning algorithm is used to compute protective perturbations, starting with forward propagation optimization to achieve rapid convergence, followed by iterative backward propagation learning to alleviate overfitting. This strategy offers state-of-the-art transferability and protection across various systems. With the running of only three steps, our Fast Preemption framework outperforms benchmark training-time, test-time, and preemptive adversarial defenses. We have also devised the first, to our knowledge, effective white-box adaptive reversion attack and demonstrate that the protection added by our defense strategy is irreversible unless the backbone model, algorithm, and settings are fully compromised. This work provides a new direction to developing proactive defenses against adversarial attacks.
1 Introduction
Deep learning technologies have been demonstrated to be sensitive to adversarial attacks [16]. An adversarial attacker can utilize this weakness to manipulate medias such as images, audios, and videos. For example, the attacker can change the identity of a face image to that of the attacker without changing the looking, enabling the attacker to illegally access the face authentication system [43]. The attacker can also add slight noise on traffic signs, possibly causing an accident during autonomous driving [13]. Defending against such attacks is a critical but unsolved challenge. Passive defenses [5, 26, 31, 20], on the one hand, trade off classification accuracy for resistance against threats [46], but may be vulnerable to counterattacks [6, 42]. Proactive defenses [29, 15], on the other hand, do not achieve satisfactory time efficiency when running protection in real-time applications. Additionally, these proactive defenses typically require paired modules for protection and inference, so their protection is untransferable to other systems. Furthermore, if a provided model is used, the user may concern if backdoor exists in this model [35].
Figure 1: A preemptive defense offers transferable protection against an adversarial attack by preemptively manipulating the image before it is attacked.
To address these problems, we devise a proactive strategy, called Fast Preemption, as illustrated in Fig. 1. Fast Preemption preemptively “attacks” (manipulates for enhancement) the medias from the user side before being attacked by the third party. Specifically, Fast Preemption learns robust features and predicts possible adversarial perturbations, then enhances these robust features and reverses the predicted perturbations for purpose of protection. In this case, when these protected medias are further attacked, these adversarial perturbations will be automatically neutralized. This strategy has three distinct advantages. (i) A proactive preemptive defense provides better resistance against threats than a passive one without degrading accuracy, whereas Fast Preemption requires no additional operation for the classification system. (ii) Fast Preemption allows plug-and-play implementation because it does not require paired modules, alleviating the security and privacy concerns through a backdoor when using another provider’s system [14, 35]. (iii) From the attacker’s perspective, once a training- or test-time defense (typically integrated within a classification model) is breached, all inputs are affected. In contrast, with this preemptive defense, only a single user is affected as users are likely to be using different settings. Therefore, Fast Preemption meets four basic requirements:
  • Preserve clean accuracy and improve robust accuracy: Basic requirements for all defenses.
  • Not require ground truths: It is impractical to label each input manually.
  • Provide an efficient defense: Time cost per process must be satisfactory for user interaction.
  • Be transferable to other systems: Effect of protection must be transferable across various systems, so users do not have to use a paired module.
We propose the Fast Preemption defense includes three components: a classifier model for automatically labeling inputs, a backbone model for learning features (differ from the classifier), and the forward-backward (Fore-Back) cascade learning algorithm for efficiently computing transferable protective perturbations. We conducted evaluation on two datasets [24, 1] across six models [17, 5, 32, 26, 29, 11], defending against two typical adversarial attacks [28, 9]. The results demonstrate the superior performance of the Fast Preemption strategy in terms of clean and robust accuracies, transferability, and time cost compared with those of benchmark training-time, test-time, and preemptive adversarial defenses [5, 26, 31, 29, 15]. For example, in an end-to-end comparison with the state-of-the-art (SOTA) preemptive defense [29] on the ImageNet dataset [1], Fast Preemption attained 1.5% and 5.6% increases in terms of the clean and robust accuracies, respectively, against AutoAttack (Standard) [9] and was 33 times faster.
Shannon insightfully claimed that “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them [39].” On the basis of that assumption, we designed the first effective white-box adaptive reversion attack, dubbed Preemptive Reversion, for evaluating the reliability of preemptive defenses, i.e., whether they are reversible. Preemptive Reversion conducts the secondary preemptive defense upon the protected example and then reverses these new perturbations because, in the same setting, the secondary defense offers mostly the same perturbations as the first defense. Along with adaptive reversion, we propose using a more comprehensive evaluation protocol, one that requires a decrease in the clean accuracy after the instance is preemptively defended (which is ensured by selecting a poor classifier). If clean accuracy increases after reversion, the reversion succeeded. If it decreases after reversion (due to distortion), reversion failed.
The results from reversion attack testing demonstrate that the Preemptive Reversion attack can smartly erase protective perturbations when all defense details are compromised (i.e., a white-box situation), whereas blind JPEG distortion and adversarial purification [31] only further distort an image. However, this white-box situation is impractical with the Fast Preemption strategy for real-world applications because the proposed framework enables users to individually customize their defense settings. Apart from a white-box situation, Fast Preemption protection cannot be reversed using existing reversion approaches.
We summarize our contributions as follows: (i) We present a new preemptive defense strategy (Fast Preemption) providing more transferable protection without requiring ground truths or requiring users to use a paired module. It is the first to use different models for the classifier and backbone. (ii) We present a forward-backward (Fore-Back) cascade learning algorithm that speeds up preemptive defense for efficient user interaction. (iii) We present the first effective adaptive reversion attack (Preemptive Reversion) for evaluating the reliability of preemptive defenses in a white-box setting and a more comprehensive protocol to evaluate whether reversion succeeds.
2 Related Work
Existing adversarial defenses can be categorized into three primary types: training-time, test-time, and preemptive defenses. Appendix A provides a more detailed review.
Training-time defenses produce robust deep learning models that may correctly classify adversarial examples (AEs) as their ground truth labels [16, 38, 37, 17, 11, 45, 12, 32, 46, 5, 26, 40]. However, they have two shortcomings. First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness. Second, training-time defenses typically reduce clean accuracy in favor of robust accuracy.
Test-time defenses are divided to detection and purification. Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs [7, 47, 25, 20]. However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks. Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy. In contrast, adversarial purification aims to purify AEs by eliminating or invalidating perturbations. Earlier studies used image distortion for purification [44, 18, 33]. Lately, more practical approaches generate benign replacements for inputs regardless of whether they are malicious [34, 31]. The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy, and the possibility of being counterattacked [6, 42].
Preemptive defenses seek to generate robust examples (REs) that replace the original images, which are subsequently discarded [36, 29, 15]. These REs are inherently more difficult to manipulate, thereby providing resistance against adversarial attacks. However, such approaches face significant limitations, including high computational costs, limited transferability across models, and, in some cases, lower robustness compared to other defense mechanisms.
Lastly, Alfarra et al. [2] proposed a test-time defense that integrates the preemptive defense technique as an anti-adversary layer. However, the use of preemptive defense techniques for test-time defenses is less practical because test-time defenses do not know whether inputs are malicious, so an anti-adversary layer may fail to resist certain attacks because it relies on a label that was predicted from the original classifier and may initially be incorrect.
Figure 2: Fast Preemption framework. The classifier labels inputs automatically, which improves clean accuracy and thereby improves defense performance. The backbone model learns crucial features, so it must be adversarially trained. The Fore-Back learning algorithm first runs forward propagation ($x_{0}\rightarrow x_{1}$) to quickly move away from the incorrect classes and then runs backward propagation ($x_{1}\rightarrow x_{2}$) to alleviate overfitting on the backbone model.
3 Preemptive Defense: Fast Preemption
With the preemptive defense strategy, the image is preemptively “attacked” (manipulated for enhancement), protecting it before being attacked. In this section, we present the framework and algorithm of our Fast Preemption strategy.
3.1 Problem Definition
Adversarial attacks and preemptive defenses both manipulate inputs by adding subtle perturbations. Specifically, let $x$ and $y$ denote the original image and its ground-truth label, respectively, and let $f_{v}$ denote the victim model. Let $x^{\prime}$ denote the manipulated example (AE or RE), formulated as
$x^{\prime}=x+\delta,\ s.t.\ ||\delta||_{p}\leq\epsilon,$ (1)
where $\delta$ denotes the perturbations, $||\cdot||_{p}$ represents the $L_{2}$- or $L_{\infty}$-norm and $\epsilon$ is the setting of the maximum distance. Therefore, adversarial attacks and preemptive defenses aim to solve different optimization problems (as explained below) by changing $\delta$ with the inputs of $x$, $y$, and $f_{v}$.
3.1.1 Adversarial Attack
An adversarial attack leads to misclassification due to subtle perturbations. Given an original image $x$ and its ground truth $y$, the adversarial attacker is able to generate an AE by maximizing the cross entropy (CE) loss of $x$ for label $y$, thereby reducing the model’s confidence in correctly classifying the input as its true label. Therefore, the optimization objective of an adversarial attack for misclassification is
[…]
3.1.2 Preemptive Defense
A preemptive defense aims to generate REs those will make Eq. 2 less successful. With the preemptive defense strategy, “backward propagation optimization” is used to reverse the adversarial attack perturbations [29]. In other words, the optimization of Eq. 2 remains, but output $\delta$ is reversed. The backward propagation optimization is formulated as
[…]
In contrast to backward propagation optimization, we newly define “forward propagation optimization” for the preemptive defense strategy. It directly reverses the optimization direction of Eq. 2 from maximization to minimization, which in turn enhances the model’s confidence in accurately classifying the input according to its true label:
[…]
3.2 Framework
As illustrated in Fig. 2, the Fast Preemption framework includes three components.
Classifier: Correct input labels determines the correct optimization direction for defense. However, for efficient user interaction, it is impractical to label each input manually. Classifier $f_{c}$ labels these inputs automatically. Therefore, $y$ in Eq. 3 and Eq. 4 can be replaced by $f_{c}(x)$. The framework is integrated with user application before being attacked, so $f_{c}$ does not have to show any transferability or resistance against threats, but it must offer the best classification accuracy. Processing images with ground truths can be regarded as using a perfect classifier.
Backbone Model: With Fast Preemption, the input image is iteratively optimized on backbone model $f_{b}$, which may differ from victim model $f_{v}$ in Eq. 3 and Eq. 4. Therefore, due to its complexity, $f_{b}$ determines the transferable clean and robust accuracies and time cost. To achieve better performance, the backbone model $f_{b}$ must be adversarially trained so it can learn more crucial features instead of random pixels. The limitations of adversarially trained models (i.e., training-time defenses), such as the degradation of clean accuracy, are alleviated by the well-performing nonrobust classifier $f_{c}$ in our framework. As a result, the preemptive defense optimizes the inputs towards the correct labels. Therefore, two different models, classifier $f_{c}$ and backbone $f_{b}$, are integrated into the framework.
Fore-Back Cascade Learning Algorithm: To meet our framework, the optimization problems formulated as Eq. 3 and Eq. 4 can be reformulated as
[…]
respectively, where $f_{c}$ is our classifier and $f_{b}$ is our backbone model. This cascade learning algorithm is described in detail in the next section.
3.3 Forward-Backward Cascade Learning
The Fore-Back cascade learning algorithm is used to compute protective perturbations. The input is first moved away from the classification boundaries of all incorrect classes so that the manipulated input will be easier to be classified as the ground truth. We define this as forward propagation ($\rightarrow$). Forward propagation can quickly attain convergence (less time cost) but causes overfitting on the backbone model so that the defense is less transferable to other victim models. Backward propagation optimization ($\leftarrow$) is thus conducted following forward propagation. Backward propagation reverses the adversarial perturbations to make them protective perturbations, which effectively alleviates overfitting and thereby improves transferability. In summary, the forward and backward propagation both enhance robustness, but in different optimization directions. The complete algorithm is presented in Algorithm 1 of Appendix B.
3.3.1 Loss Function
From Eq. 5 and Eq. 6, the loss function of our Fore-Back cascade learning algorithm is the CE loss:
$\ell(x_{i}|x,f_{c},f_{b})=CE(f_{b}(x_{i}),f_{c}(x)),$ (7)
where $x_{i}$ and $x$ denote the output example of the $i$-th iteration and the original input, respectively, and $f_{c}$ and $f_{b}$ are our classifier and backbone models, respectively. The forward optimization reduces $\ell$, while the backward optimization increases it.
3.3.2 Preliminary - Backward Propagation
The use of backward propagation for preemptive defense was introduced in the SOTA preemptive defense strategy [29]. At each step of that strategy, a projected gradient descent (PGD) attack [28] is run first as $x_{1}\rightarrow x_{1}^{a}$ (Fig. 2), and then the adversarial perturbations are reversed in accordance with the learning rate and thereby forms protective perturbations, as $x_{1}\rightarrow x_{2}$.
Specifically, backward propagation first runs an $I$-iteration $L_{\infty}$-norm PGD attack from $x_{t}$, which is the output of the previous optimization step (forward or backward). In other words, the backward propagation increases $\ell$ via
$x_{t,0}=x_{t},$ (8)
$x_{t,i+1}=\prod_{x,\epsilon}(x_{t,i}+\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
$s.t.\ ||x_{t,i+1}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
where $\epsilon$ is the setting of the maximum distance and $\alpha$ is the step size. The final output of Eq. 8 is AE $x_{t}^{a}$. Next, backward propagation subtracts perturbations $x_{t}^{a}-x_{t}$ from $x_{t}$ in accordance with learning rate $\eta$, thereby computing the next robust example, $x_{t+1}$:
$x_{t+1}$ $=\prod_{x,\epsilon}(x_{t}-\eta\cdot(x_{t}^{a}-x_{t}))$ (9)
$=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))).$
The SOTA preemptive defense strategy [29] conducts only backward propagation learning. However, adversarial attacks can be randomly initialized, so backward propagation requires many steps to converge (minimum ten steps). The Fast Preemption strategy accelerates defense by running forward propagation before running backward propagation learning.
3.3.3 Combining Forward and Backward Propagation
We define forward propagation learning as a direct enhancement by moving away from the decision boundary in accordance with the learning rate ($x_{0}\rightarrow x_{1}$ in Fig. 2). Therefore, it first reduces $\ell$ via
$x_{t,0}=x_{t},$ (10)
$x_{t,i+1}=\prod_{x,\epsilon}(x_{t,i}-\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
$s.t.\ ||x_{t,i+1}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
the final output of which is robust example $x_{t}^{r}$. Subsequently, protective perturbations $x_{t}^{r}-x_{t}$ are directly added to $x_{t}$ in accordance with learning rate $\eta$:
$x_{t+1}$ $=\prod_{x,\epsilon}(x_{t}+\eta\cdot(x_{t}^{r}-x_{t}))$ (11)
$=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
which is the same as for Eq. 9. Therefore, the above derivations can be summarized as
$x_{t,0}=x_{t},$ (12)
$x_{t,i+1}=\left\{\begin{array}[]{rl}\prod_{x,\epsilon}(x_{t,i}-\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b})))&\rightarrow\\ \prod_{x,\epsilon}(x_{t,i}+\alpha\cdot sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b})))&\leftarrow\end{array}\right.$
$x_{t+1}=\prod_{x,\epsilon}(x_{t}-\eta\cdot\alpha\cdot\sum\limits_{i=1}^{I}sign(\nabla\ell(x_{t,i}|x,f_{c},f_{b}))),$
$s.t.\ ||x_{*}-x||_{\infty}\leq\epsilon\And 0\leq i<I-1,$
where $x_{*}$ denotes $x_{t,i+1}$ or $x_{t+1}$. Equation 12 indicates that both forward and backward propagation make inputs more robust but with different learning directions for loss $\ell$.
Whereas backward propagation requires many steps to converge, forward propagation converges quickly. However, forward propagation causes overfitting of the backbone model, so the defense will be less transferable to other victim models. Fast Preemption alleviates this overfitting: $T_{\rightarrow}$ steps of forward propagation are run until there is slight overfitting; then $T_{\leftarrow}$ steps of backward propagation are run to eliminate the overfitting. We evaluated whether the learning is overfitted by evaluating transferability. With this combination of forward and backward propagation, Fast Preemption attains SOTA performance with only three steps (one step $T_{\rightarrow}$, followed by two steps $T_{\leftarrow}$), which dramatically reduces the time cost of RE generation.
3.3.4 Averaging $N$ Examples
Forward and backward propagation both require a backbone optimization algorithm, i.e., an adversarial attack algorithm. For example, The Bi-Level optimization [29] used PGD [28] for backward propagation. However, PGD has to run many sequential iterations to compute superior adversarial perturbations, which greatly increases the time cost. With the Fast Preemption strategy, Fast Gradient Sign Method (FGSM) [16] is used to speed up the computation of adversarial perturbations.
Since FGSM has difficulty obtaining optimal perturbations, with the Fast Preemption strategy, it is run $N$ times with random initialization to produce $N$ FGSM examples, The perturbations are then averaged to obtain the final output. We demonstrate that averaging the perturbations for PGD examples negligibly affects performance, whereas with the Fast Preemption strategy, FGSM optimization is effectively improved while efficiency is preserved.
Furthermore, averaging perturbations in the $N$ FGSM examples is faster than processing a single example with iterative PGD. To generate a single RE, iterative PGD has to run each iteration sequentially from a single input, whereas $N$ FGSM examples can be produced by repeating the single input $N$ times and processing the results in parallel.
4 Adaptive Attack: Preemptive Reversion
A reversion attack is used to evaluate whether a preemptive defense is reversible. In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks. Preemptive Reversion is the only feasible approach to reverse preemptive defense in the white-box scenario. Nevertheless, without full knowledge of the backbone model and settings, our Fast Preemption defense is hard to reverse, so it is reliable.
4.1 Algorithm
When running preemptive defenses sequentially for the same input and settings, the perturbations are mostly the same for two rounds (see Fig. 5). On the basis of this finding, we devised the first effective white-box adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Fig. 3. The complete algorithm is presented in Algorithm 2 of Appendix B. Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$. Then, the Preemptive Reversion algorithm reverses the secondary perturbations from $x^{r}$ to obtain $\tilde{x}$, which is the prediction of the original input $x$:
$\tilde{x}=x^{r}-((x^{r})^{r}-x^{r}).$ (13)
Figure 3: Illustration of Preemptive Reversion algorithm. It first conducts a secondary preemptive defense on RE $x^{r}$ and then reverses the perturbations of the secondary defense: $(x^{r})^{r}-x^{r}$.
4.2 Attack Scenarios
We define two scenarios for a reversion attack in accordance with the knowledge the attacker has about the backbone model and defense algorithm:
  • White-Box: The attacker knows both the backbone model and defense algorithm (including settings), so the attacker can conduct a secondary defense the same as the user does. This is the strongest, yet impractical, assumption because different users may use different settings.
  • Black-Box: The attacker knows the defense algorithm (typically public) but not the backbone model (infinite selections for users). The challenge for the attacker in this scenario is that the crucial features learned in the secondary defense may differ from those learned by the user.
Our proposed Preemptive Reversion attack is the first effective reversion attack for evaluating the reversibility of preemptive defenses, including our Fast Preemption defense, but it is feasible only in the white-box scenario. Without full knowledge of the defense, Fast Preemption is reliable.
4.3 Evaluation Protocol
If a reversion attack succeeds, the improved clean accuracy should drop. However, image distortion can create a similar effect. Consequently, a drop in clean accuracy does not necessarily indicate the success of a reversion attack. To address this issue, we proposed a comprehensive protocol that can distinguish successful reversion from distortion. First, we purposely give incorrect labels to partial inputs to guarantee that the clean accuracy after being defended is lower than the original accuracy. In this situation, if the clean accuracy approaches close to the original accuracy, the reversion attack succeeded. Otherwise, the reversion attack failed, and the deterioration was caused by distortion.
5 Experiments
5.1 Datasets, Classifier, and Backbone Models
We conducted experiments on two datasets: CIFAR-10 [24] and ImageNet [1]. For CIFAR-10, ResNeXt29-32x4d [19] and PreActResNet-18 [17] were selected as the classifier and backbone model, respectively. For ImageNet, DeiT Base [41] and ConvNeXt-L+ConvStem [40] were selected as the classifier and backbone model, respectively. The model architectures are listed in Tab. 3 of Appendix C.
5.2 Attacks and Defenses
Adversarial attacks. We evaluated defenses against two adversarial attacks: PGD [28] and AutoAttack [9]. For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, and FAB [8] and Square Attack [3]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [4]). Reversion attacks. We ran reversion attacks to evaluate: (i) whether the protected samples induced by preemptive defenses cannot be easily and precisely reversed back to their original unprotected status, which we define as “irreversible” in this paper; (ii) whether the proposed adaptive reversion attack, Preemptive Reversion, is effective. Therefore, apart from the specific Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [31]).
Training-time defenses: We compared Fast Preemption with WideResNet-94-16 [5] and Swin-L [26] on the CIFAR-10 and ImageNet datasets, respectively. The model architectures of these defenses differ from the backbone models. Test-time defenses: We used DiffPure [31] as the benchmark test-time defense. The inputs were purified by DiffPure and then fed into WideResNet-94-16 [5] or Swin-L [26] in accordance with the dataset used. Preemptive defenses. We compared Fast Preemption with the SOTA preemptive defense, Bi-Level [29], and the latest preemptive defense, $A^{5}$ [15].
5.3 Evaluation metrics
We used clean accuracy and robust accuracy for performance evaluation, where higher accuracy is better from the defense perspective. Here, accuracy refers to the ratio of predictions that match their ground-truth labels. Clean accuracy was computed for benign images (original and REs) while the robust accuracy was computed for AEs.
Table 1: End-to-end performance against AutoAttack [9] and comparison with benchmark adversarial defenses.
Dataset Defense Time Cost (s) Clean (%) Standard $L_{\infty}$ (%) Standard $L_{2}$ (%) Random $L_{\infty}$ (%)
CIFAR-10 Training-Time WideResNet-94-16 [5] 0 93.4 74.6 69.6 74.8
8
Test-Time DiffPure [31] 0.3 91.4 78.1 84.9 78.4
8
Preemptive $A^{5}$ [15] 0.01 93.5 73.8 69.0 74.3
Bi-Level [29] 28 92.5 81.4 78.6 81.6
Fast Preemption (ours) 0.04 95.0 86.5 84.3 86.7
8
Fast Preemption (ours) + DiffPure [31] 0.34 94.2 89.1 92.8 89.7
ImageNet Training-Time Swin-L [26] 0 93.0 67.0 59.5 67.5
8
Test-Time DiffPure [31] 1.2 88.8 73.7 79.4 73.8
8
Preemptive Bi-Level [29] 43 93.6 71.5 62.8 72.3
Fast Preemption (ours) 1.3 95.1 77.1 67.4 78
8
Fast Preemption (ours) + DiffPure [31] 2.5 93.3 82.9 87.6 83.4
  • Bold indicates the best performance.
6 Performance
In this section, we present the end-to-end evaluation results and comparisons with benchmarks. The results were obtained in the settings listed in Tab. 4 of Appendix D. The time cost was recorded with an NVIDIA A100 40GB GPU, as described in Appendix E.
Table 2: Preemptive defenses resist reversion attacks.
Original (%) Preemptive Defense Distortion Purification Preemptive Reversion (ours)
Classifier Backbone Approach Clean (%) JPEG (%) DiffPure (%) White (%) Black (%)
93.5 Bi-Level Bi-Level Bi-Level 91.9 90.5 90.8 93.1 90.7
Bi-Level PreActResNet-18 Bi-Level 91.2 90.6 90.7 93.4 91.0
Bi-Level PreActResNet-18 Ours 91.3 90.5 90.5 93.4 90.6
ResNeXt29-32x4d PreActResNet-18 Ours 95.5 94.4 94.8 94.0 94.6
  • Numbers in red demonstrate that only our reversion attack is feasible but only in the white-box setting.
Figure 4: Perturbations of adversarial attacks and proposed preemptive defenses. Observation of differences between original image and RE show that Fast Preemption perturbations are unnoticeable. Comparison of perturbations between AutoAttack and Fast Preemption show that preemptive defense is more likely to succeed when perturbations are similar.
6.1 Robustness against Adversarial Attacks
We evaluated transferable defense performance for the situation in which the defense has no knowledge of the AutoAttack [9] and victim models [5, 26]. Table 1 compares the performance of Fast Preemption with that of the benchmark defenses against AutoAttack [9]. Fast Preemption performance was better than that of the benchmarks in terms of clean and robust accuracies, whereas its time cost was competitive. For example, in comparison with the least time costly method WideResNet-94-16 [5] with the CIFAR-10 dataset, Fast Preemption had 1.6% and 11.9% better clean and robust accuracies (Standard $L_{\infty}$), respectively, with 0.04 more seconds per generation. Compared with the SOTA preemptive defense, Bi-Level [29], which had the second best accuracy, Fast Preemption had 2.5% and 5.1% better clean and robust accuracies (Standard $L_{\infty}$), respectively and was 700 times faster. Note that preemptive defenses can work together with training-time and test-time defenses. For example, when combined with a test-time defense, the performance of Fast Preemption was remarkably improved, as reflected in the last rows of CIFAR-10 and ImageNet in Tab. 1, respectively.
Figure 4 illustrates the results of qualitative analysis. A visual comparison of the original image and RE revealed that the perturbations generated by Fast Preemption are unnoticeable. Moreover, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations are similar to the attack ones. Note that the perturbation illustration $\delta_{Gray}$ was drawn by first normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$ to produce $\delta_{RGB}$ and then transforming $\delta_{RGB}$ into grayscale using the National Television Standards Committee (NTSC) formula [27]:
$\delta_{RGB}=$ $\ \frac{\delta+\epsilon}{2\cdot\epsilon},$ (14)
$\delta_{Gray}=$ $\ 0.299\cdot\delta_{RGB}.R+$
$\ 0.587\cdot\delta_{RGB}.G+0.114\cdot\delta_{RGB}.B,$
where $\delta_{RGB}.R$, $\delta_{RGB}.G$, and $\delta_{RGB}.B$ represent the red, green, and blue channels, respectively.
6.2 Resistance against Reversion Attacks
We conducted reversion attacks to evaluate whether preemptive defenses are reversible. Apart from the proposed Preemptive Reversion attack, we used JPEG compression and adversarial purification because they require no knowledge of the defense. Preemptive Reversion was evaluated in both white- and black-box scenarios (see Sec. 4.2). The results are summarized in Tab. 2. The results shown in red approach the original accuracy, demonstrating that only our reversion attack was able to reverse the preemptive defense. The other reversion attacks failed because they further reduced clean accuracy (due to distortion). Nevertheless, Preemptive Reversion is only feasible in the white-box setting. Without full knowledge of the defense, Fast Preemption is irreversible. Figure 5 further supports our claim that the white-box reversion successfully erases most protective perturbations while the black-box reversion fails. However, the white-box reversion cannot fully remove the protective perturbations due to the randomization in Fast Preemption (i.e., randomly initialized FGSM).
Figure 5: Perturbations before and after proposed Preemptive Reversion attack. White-box reversion successfully erased most protective perturbations while black-box reversion failed.
Moreover, the results in the fourth row of Tab. 2, which were obtained without using our evaluation protocol, demonstrate that our protocol for evaluating reversion attacks is more comprehensive. The results shown in the other rows demonstrate that our Preemptive Reversion in the white-box setting is the only feasible reversion attack as it is the only one for which clean accuracy was higher after reversion. However, the lower numbers in the fourth row indicate that the decrease in accuracy with Preemptive Reversion in the white-box setting was due to successful reversion, whereas the other decreases were due to distortion. Therefore, if only the results in the fourth row are considered, a decrease in accuracy does not necessarily indicate the effectiveness of a reversion attack.
7 Ablation Studies
We conducted ablation studies on the Fast Preemption framework and its learning algorithm using the CIFAR-10 dataset. We used AutoAttack (Standard) [9] against a black-box target model RaWideResNet-70-16 [32], employing $L_{\infty}$-norm perturbations with $\epsilon=8/255$. As demonstrated in Fig. 6, Fast Preemption significantly outperforms the SOTA Bi-Level method [29], increasing clean accuracy from 92% to 95.3%, improving robust accuracy from 80.1% to 85.6%, and reducing computation time from 28 seconds to just 0.04 seconds. These studies were conducted to identify the optimal configurations, with more extensive ablation results presented in Appendix F.
Figure 6: Compared with the SOTA preemptive defense Bi-Level [29], our framework improved the clean and robust accuracies, and our learning algorithm reduced the time cost for defense.
8 Conclusion
We have presented an efficient transferable preemptive adversarial defense that we call Fast Preemption. We demonstrated that with its framework and Fore-Back cascade learning algorithm, Fast Preemption offers transferable protection against adversarial attacks among unknown victim models without requiring ground truths. In terms of the time cost, clean accuracy, and robust accuracy, Fast Preemption outperformed benchmark training-time, test-time, and preemptive defenses. We also demonstrated that the protection from Fast Preemption is irreversible unless the backbone model, algorithm, and settings of the defense are all compromised. This paper lays the theoretical foundation for a preemptive defense framework that can shape future research on each component’s optimization instead of end-to-end application in the relevant field.
Acknowledgments
This work was supported by JSPS KAKENHI Grants JP21H04907 and JP24H00732, by JST CREST Grants JPMJCR18A6 and JPMJCR20D3 including AIP challenge program, by JST AIP Acceleration Grant JPMJCR24U3, by JST K Program Grant JPMJKP24C2 Japan, and by the project for the development and demonstration of countermeasures against disinformation and misinformation on the Internet with the Ministry of Internal Affairs and Communications of Japan.
Supplementary Material
This supplementary material provides an in-depth discussion of individual related works (Appendix A), the algorithms underpinning the proposed adversarial defense and adaptive reversion attack (Appendix B), model architectures (Appendix C), and the experimental configurations for both attacks and defenses (Appendix D). Additionally, it details the implementation specifics (Appendix E) and presents comprehensive ablation study results Appendix F, covering aspects such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples.
Appendix A Related Work
Existing adversarial defenses are divided into three categories: training-time, test-time, and preemptive.
Training-time defenses produce robust deep learning models that may correctly classify adversarial examples (AEs) as their ground truth labels. The earliest mechanisms, e.g., adversarial training [16], incorporated various types of AEs along with their ground-truth labels in the training dataset. Zhang et al. [46] proposed TRADES, which introduces a theoretically grounded approach to balance robustness and accuracy in adversarial training by incorporating a custom loss function that controls the trade-off between clean accuracy and adversarial robustness. To improve the efficiency of training on large-scale datasets, Shafahi et al. [38] devised accelerated adversarial training in which backward-pass computations are reused. This entails curating a large number of specific real-world samples; Sehwag et al. [37] circumvented this problem by using additional data from proxy distributions learned by advanced generative models for adversarial training. Gowal et al. [17] demonstrated that generative models trained solely on the original dataset can be leveraged to artificially increase the training dataset’s size and improve adversarial robustness against $L_{p}$-norm perturbations. Debenedetti et al. [11] proposed using a custom recipe of vision transformers for adversarial training in which weaker data augmentation is used with warm-up in the augmentation intensity. Inspired by denoising diffusion models [21], Wang et al. [45] conducted adversarial training using data generated by an elucidating diffusion model [22]. Dong et al. [12] introduced robust examples (REs) generated by a preemptive defense into the adversarial training and demonstrated alleviation of clean accuracy deterioration. Peng et al. [32] devised a robust model which follows a suite of generalizable robust architectural design principles: (i) an optimal range for depth and width configurations; (ii) a preference for a convolutional rather than patchify stem stage, and (iii) a robust residual block design featuring squeeze and excitation blocks and nonparametric, smooth activation functions. Bartoldson et al. [5] developed the current state-of-the-art (SOTA) robust model by scaling laws for adversarial training to assess robustness in image classifiers, revealing that even with significant computational resources, robustness asymptotes below human-level performance due to flaws in attack formulations generating invalid data, suggesting future efforts should address both computational efficiency and image validity.
Training-time defenses have two shortcomings. First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness. Second, training-time defenses typically reduce clean accuracy in favor of robust accuracy, even the one proposed by Dong et al. [12], which was designed to circumvent this problem.
Test-time defenses are divided to detection and purification. Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs. Among recent studies, Cohen et al. [7] devised a k-nearest neighbor model for ranking the most supportive training samples for any given validation example in training a detector. Zuo and Zeng [47] proposed using a binary classifier, erase-and-restore, that uses the intriguing sensitivity of $L_{2}$-norm attacks by randomly erasing certain pixels in an $L_{2}$-norm AE and then restoring them with an inpainting technique. Liang et al. [25] detected AEs by comparing the classification results of a given sample and a denoised version that was generated through adaptive noise reduction combining scalar quantization and a smoothing spatial filter. Most recently, Hickling et al. [20] devised a detector using explainable deep reinforcement learning for uncrewed aerial vehicle agents. However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks [6]. Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy.
In contrast, adversarial purification aims to purify AEs by eliminating or invalidating perturbations. Earlier studies used image distortion for purification. For instance, Wang et al. [44] distorted images by randomly nullifying pixels. Guo et al. [18] preprocessed inputs with a nondifferentiable transformation. Raff et al. [33] combined a large number of individually weak defenses into a single randomized transformation to build a strong defense. Lately, more practical approaches generate benign replacements for inputs regardless of whether they are malicious. Ren et al. [34] trained a generative model using autoencoder technology [23]. They claimed the existence of an immune space, in which perturbations have fewer adverse effects, and they inactivated adversarial perturbations by restricting them to that subspace. Nie et al. [31] devised the DiffPure benchmark, which uses denoising diffusion models [21]. DiffPure replaces perturbations with Gaussian noise and then restores the clean image by reversing the diffusion process. The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy, and the possibility of being counterattacked [42].
Lastly, Alfarra et al. [2] proposed a test-time defense that integrates the preemptive defense technique as an anti-adversary layer. However, we believe that the use of preemptive defense techniques for test-time defenses is less practical because test-time defenses do not know whether inputs are malicious, so an anti-adversary layer may fail to resist certain attacks because it relies on a label that was predicted from the original classifier and may initially be incorrect.
Algorithm 1 Fast Preemption
1:Original image $x$, classifier $f_{c}$, backbone model $f_{b}$, maximum $L_{\infty}$-norm distance $\epsilon$, learning rate $\eta$, maximum forward steps $T_{\rightarrow}$ and backward steps $T_{\leftarrow}$, number of examples for averaging perturbations $N$.
2:Robust example $x^{r}$.
3:// Forward Propagation
4:$x_{0}=x$
5:for $t<T_{\rightarrow}$ do
6:       for $n=1,...,N$ do
7:             $\delta_{n}\sim U(-\epsilon,+\epsilon)$$\triangleright$ random initialization
8:             $x_{t,n}=x_{t}+\delta_{n}$
9:             $x_{t,n}^{r}=\prod_{x,\epsilon}(x_{t,n}-\epsilon\cdot sign(\nabla\ell(x_{t,n}|x,f_{c},f_{b})))$$\triangleright$ FGSM decreases $\ell$
10:       end for
11:       $\delta_{t}^{r}=\frac{1}{N}\sum\limits_{n=1}^{N}x_{t,n}^{r}-x_{t}$$\triangleright$ averages perturbations
12:       $x_{t+1}=x_{t}+\eta\cdot\delta_{t}^{r}$$\triangleright$ direct enhancement
13:end for
14:// Backward Propagation
15:$x_{0}=x_{T_{\rightarrow}}$
16:for $t<T_{\leftarrow}$ do
17:       for $n=1,...,N$ do
18:             $\delta_{n}\sim U(-\epsilon,+\epsilon)$$\triangleright$ random initialization
19:             $x_{t,n}=x_{t}+\delta_{n}$
20:             $x_{t,n}^{a}=\prod_{x,\epsilon}(x_{t,n}+\epsilon\cdot sign(\nabla\ell(x_{t,n}|x,f_{c},f_{b})))$$\triangleright$ FGSM increases $\ell$
21:       end for
22:       $\delta_{t}^{a}=\frac{1}{N}\sum\limits_{n=1}^{N}x_{t,n}^{a}-x_{t}$$\triangleright$ averages perturbations
23:       $x_{t+1}=x_{t}-\eta\cdot\delta_{t}^{a}$$\triangleright$ reverses the attack
24:end for
25:return $x^{r}=x_{T_{\leftarrow}}$
Algorithm 2 Preemptive Reversion
1:Robust example $x^{r}$.
2:Reversed example $\tilde{x}$.
3:$(x^{r})^{r}=PreemptiveDefense(x^{r})$$\triangleright$ secondary defense
4:$\tilde{x}=x^{r}-((x^{r})^{r}-x^{r})$$\triangleright$ reverses the secondary defense
5:return $\tilde{x}$
Preemptive defenses aim to generate REs that replace original images, which are then discarded. These REs are more challenging to manipulate and thus resist adversarial attacks. Salman et al. [36] was the first to devise a preemptive defense strategy utilizing REs, which they referred to as “unadversarial” examples. They were learned using projected gradient descent (PGD) [28] and improved both in-distribution performance and robustness against unforeseen perturbations. Moon et al. [29] devised a Bi-Level optimization algorithm (abbreviated as Bi-Level in this paper) that finds points in the vicinity of natural images that are robust against adversarial perturbations; it achieved SOTA robust accuracy. However, its time cost is huge, and its transferability is low. Frosio and Kautz [15] introduced the $A^{5}$ (“Adversarial Augmentation Against Adversarial Attacks”) framework that uses an individual classifier and a robustifier (a generator), which are co-trained. It is much more efficient than the Bi-Level optimization algorithm, but it is not as robust.
Appendix B Algorithms of Fast Preemption and
Preemptive Reversion
The complete defense algorithm of Fast Preemption is presented in Algorithm 1. With PyTorch implementation, Steps 4 to 8 and Steps 15 to 19 can be respectively processed in parallel. The complete adaptive attack algorithm of Preemptive Reversion is presented in Algorithm 2.
Appendix C Model Architectures
We evaluated attacks and defenses using models with various architectures, as listed in Tab. 3.
Table 3: Model architectures.
      Dataset       Architecture
      CIFAR-10       WideResNet-94-16 [5]
      RaWideResNet-70-16 [32]
      WideResNet-34-20 [29]
      XCiT-L12 [11]
      PreActResNet-18 [17] (backbone)
      ResNeXt29-32x4d [19] (classifier)
      ImageNet       Swin-L [26]
      ConvNeXt-L+ConvStem [40] (backbone)
      DeiT Base [41] (classifier)
Table 4: Experiment Settings.
        Type         Approach         Settings
        Adversarial         PGD         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$ or $L_{2}\ \epsilon=0.5^{1}/2^{2}$
        Attack         AutoAttack
                 Reversion         Attack         JPEG         Quality Factor = 50%
        DiffPure         Diffusion Timestep = $0.1^{1}/0.15^{2}$
        Fast Preemption (ours)         Same with our defense
        Defense         DiffPure         Diffusion Timestep = $0.1^{1}/0.15^{2}$
        Bi-Level         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$, $\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
        $A^{5}$         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
        Fast Preemption (ours)         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$,$\eta^{3}=0.7$, $T_{\rightarrow}=1$, $T_{\leftarrow}=2$, $N=20$
  • 1 for CIFAR-10; 2 for ImageNet.
  • 3 $\eta$ must meet $\eta\cdot(T_{\rightarrow}+T_{\leftarrow})>1$.
Table 5: Performance with different classifiers.
       Ablation        Time Cost(s)        Transferable Defense
       Classifier        Accuracy(%)        Clean(%)        PGD(%)        AutoAttack(%)
       Ground Truths        100        28        97.2        86.9        83.1
       ResNeXt        95.9        28        95.4        85.7        82.7
       Bi-Level        89.1        29        91.9        83.9        81.0
  • To evaluate the performance of diverse classifiers in our framework, we used the same backbone model and 100 steps backward propagation as Bi-Level [29]. The transferable defenses were all evaluated using RaWideResNet-70-16 [32] as the victim model. Bold indicates the best performance.
Appendix D Experiment Settings
The full experimental settings used for evaluations are listed in Tab. 4.
Appendix E Implementation
Our implementation partially leveraged the ART [30] and RobustBench [10] leaderboards by using PyTorch. The experimental machine was equipped with an NVIDIA A100 40GB GPU.
Appendix F Ablation Studies
We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset. The attacks were PGD [28] and AutoAttack (Standard) [9] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$. As illustrated in Fig. 6 of Sec. 7, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [29] in terms of clean and robust accuracies and time cost. These studies and results were aimed at identifying the optimal strategies.
Figure 7: Different backbone models learn different features. PreActResNet-18 [17], WideResNet-94-16 [5], and Bi-Level [29] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
F.1 Classifier
In all proactive defenses, correct labeling guarantees that optimization moves in the correct direction. As demonstrated in Tab. 5, a classifier with higher accuracy leads to better performance. As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning. Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction. Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
Table 6: Performance of different backbone models.
Backbone Model: Bi-Level [29] WideResNet-94-16 [5] PreActResNet-18 [17]
Backbone Robustness Clean (%) 89.1 93.9 88.1
PGD (%) 43.9 77.1 61.8
AutoAttack (%) 42.4 74.0 58.6
Time Cost (s) for Defense 28 69 26
White-Box Defense Clean (%) 95.5 95.8 93.8
PGD (%) 84.9 91.6 85.9
AutoAttack (%) 81.8 90.2 83.3
Transferable Defense Victim: RaWideResNet-70-16 [32] Clean (%) 95.4 95.5 95.3
PGD (%) 85.7 89.7 87.6
AutoAttack (%) 82.7 88.0 85.9
Transferable Defense Victim: XCiT-L12 [11] Clean (%) 95.3 95.3 95.4
PGD (%) 72.7 78.6 79.6
AutoAttack (%) 70.8 75.7 77.3
  • To evaluate the performance of the backbone models in our framework, we followed the results from the previous section and used ResNeXt29-32x4d [19] as the classifier model, and 100 steps backward propagation as Bi-Level [29]. The transferable defenses were all evaluated using RaWideResNet-70-16 [32] and XCiT-L12 [11] as the victim models, different from the backbone models. Bold indicates the best performance.
F.2 Backbone Model
The backbone model determines the defense time cost and performance. As illustrated in Fig. 7, different backbone models learn different features. To evaluate the effect of the backbone model, we also evaluated white-box defenses, for which the backbone model is the same as the victim model, i.e., paired modules, as stated in Sec. 1, although we believe that this setting is not secure due to potential privacy leakage due to the existence of a backdoor [35].
As shown in Tab. 6, the best adversarially trained model was WideResNet-94-16 [5], so it is not surprising that its white-box defense outperformed that of the other backbone models. However, this most robust model failed to guarantee the best transferability (see results against XCiT-L12 [11]), and the time cost was much higher due to its complexity. We thus empirically selected PreActResNet-18 [17] as the backbone model in Fast Preemption as it had the least time cost and comparable transferability compared with the SOTA robust model [5]. Nevertheless, it is worth investigating more comprehensive approaches for determining the backbone model in future work.
F.3 Forward-Backward Cascade Learning
As observed in Fig. 8, forward (F) propagation accelerated defense but led to overfitting (e.g., F100-B0). In contrast, backward (B) propagation required more steps to converge (e.g., F0-B100) but alleviated overfitting by the forward propagation (e.g., F10-B90). Additionally, as shown by the curve for “Alternate,” the number of steps in forward propagation must be smaller than that in backward propagation to alleviate overfitting.
Our goal is to reduce the time cost, i.e., learning steps, so we further evaluated forward and backward propagation by limiting the learning steps, as illustrated in Fig. 9. The results show that when the number of learning steps is small, the combination of forward and backward propagation outperforms backward propagation alone (e.g., F1-B2 vs. F0-B3). Therefore, Fast Preemption runs one step forward and two steps backward propagation learning as the optimal strategy.
Figure 8: Forward propagation (F) converges faster while backward propagation (B) alleviates overfitting. The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively. The numbers in brackets represent robust accuracy against AutoAttack (Standard). “Alternate” represents the running of forward and backward propagation alternatively for 100 steps in total.
Figure 9: By combining forward (F) and backward (B) propagation, Fast Preemption obtains performance comparable to that of the SOTA preemptive defense [29] by using only three steps (F1–B2). The numbers following F and B denote the number of learning steps $T_{\rightarrow}$ and $T_{\leftarrow}$, respectively. The numbers in brackets represent robust accuracy against AutoAttack (Standard).
Table 7: Performance for different $N$ and optimization algorithms.
       Ablation        Time Cost (s)        Transferable Defense
       Optimization        $N$ Samples        Clean (%)        PGD (%)        AutoAttack (%)
       PGD        1        0.7        95.3        87.5        85.8
       FGSM        1        0.04        95.4        86.8        85.4
       PGD        20        0.8        95.2        87.7        85.9
       FGSM        20        0.04        95.5        87.8        85.9
  • Forward propagation ran one step and backward propagation ran two steps. Transferable defense was evaluated using RaWideResNet-70-16 [32] as the victim model. Bold indicates the best performance.
F.4 Optimization Algorithm
The results in Tab. 7 indicate that averaging $N$ FGSM examples results in the best performance with the least time cost. Using PGD results in much more time being spent for preemptive defense, whereas using FGSM but without averaging leads to poor performance.